OpenStudy (anonymous):

(–3)(–6)(–5) = try and SOLVE this! >:D

OpenStudy (saifoo.khan):

\[- \times - \times - = -\]

OpenStudy (tansheet):

(-3)(-6)(-5)=-90

OpenStudy (saifoo.khan):

Whereas, \[3 \times 6 \times 5 = something\]

OpenStudy (anonymous):

\[\LARGE (-)\cdot (-)\cdot (-)=(-)\] this is what @saifoo.khan wtore :) (no offense )

OpenStudy (anonymous):

wrote

OpenStudy (anonymous):

-90

OpenStudy (anonymous):

i CANT THANK U ENOUGH! ILL TRY AND DO SOMETHIN FOR YALL!

OpenStudy (saifoo.khan):

lol, my equation was looking awkward.